OpenStudy (anonymous):

Graph 4x2 + 4y2 = 64. What are the domain and range? Domain: –4 ≤ x ≤ 4 Range: –4 ≤ y ≤ 4 Domain: all real numbers Range: –16 ≤ y ≤ 16 Domain: –16 ≤ x ≤ 16 Range: –16 ≤ y ≤ 16: Domain: all real numbers Range: –4 ≤ y ≤ 4

OpenStudy (anonymous):



OpenStudy (anonymous):

gib medal

OpenStudy (anonymous):

\[4x^2 + 4y^2 = 64\] \[x^2+y^2=16\] is a circle with center at the origin, radius of 4

OpenStudy (anonymous):

Graph 4x2 + y2 = 9. What are its lines of symmetry? (1 point)